OpenStudy (jaredstone4):

Given F'(x) = f(x), and the integral from 10 to 20 of f(x)dx = 30 with F(10) = 22, find F(20).

OpenStudy (campbell_st):

well using some knowledge of integrals \[\int\limits_{10}^{20} f(x) dx = F(20) - F(10)\] so 30 = F(20) - F(10) just substitute and solve for F(20)

OpenStudy (jaredstone4):

Ohh yeah! I get it now... first Fundamental Theorem of Calculus. It was just asked in a different format and tripped me up. Thank you!